25 tonnes of mixture, has 18% H2O. How many gallons of H2O, to make mixture 50% H2O?
Answer by richwmiller(17219)

25 tonnes of mixture, has 18% H2O. How many gallons of H2O, to make mixture 50% H2O?
.18*25+x=.5*(25+x)
4.5+x=12.5+.5x
.5x=8
x=16
check
.18*25+16=.5(41)
4.5+16=20.5
ok